### v1

# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

Abstract
Commodity price volatility creates economic challenges, necessitating accurate multi-horizon forecasting. Predicting prices for commodities like copper and crude oil is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical, etc.). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity price prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. Crucially, L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings indicate the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a key advantage over traditional black-box approaches.
1 Introduction
Commodity markets, and their associated futures contracts, form a critical part of the global economy, providing essential raw materials and hedging instruments. The prices of key commodity futures, such as those for industrial metals like Copper and energy benchmarks like Crude Oil, are not merely transactional figures; they are vital indicators of economic health, drivers of industrial costs, and significant factors influencing inflation, trade balances, and geopolitical dynamics. Copper futures, for instance, are closely watched for signals about global industrial demand. Crude Oil futures are pivotal for managing energy price risk and reflect complex geopolitical and supply/demand interactions.
The inherent volatility of these commodity futures prices presents substantial risks and opportunities for a wide spectrum of market participants. Producers and consumers use futures to hedge price uncertainty, while financial institutions and traders engage in speculative and arbitrage activities. Consequently, the ability to accurately forecast commodity futures prices across various time horizons is of paramount importance for effective risk mitigation, optimized resource allocation, and informed investment decisions.
The challenge lies in the sheer complexity of factors driving futures prices. These markets are influenced by a confluence of interconnected forces:
  1. 1.
    Macroeconomic Environment: Global GDP growth, industrial production, interest rates, currency exchange rates (particularly the US Dollar), and inflation expectations.
  2. 2.
    Underlying Physical Market Fundamentals: For Copper, this includes mine production, scrap availability, and inventory levels. For Crude Oil, factors include OPEC+ decisions, non-OPEC output, geopolitical events in producing regions, and inventory levels.
  3. 3.
    Demand Fundamentals: Linked to manufacturing activity, construction, transportation needs, and energy transition dynamics.
  4. 4.
    Market Sentiment and Financialization: Speculative positioning in futures and options markets (e.g., CFTC Commitment of Traders reports), algorithmic trading, and broader risk-on/risk-off sentiment.
  5. 5.
    Other factors: Including weather patterns impacting supply/demand and inter-commodity relationships.
This intricate web of high-dimensional, often collinear, and dynamically interacting drivers makes futures price forecasting an exceptionally challenging task, characterized by non-linear relationships and regime shifts. Existing approaches often face a trade-off between predictive accuracy and interpretability. Classical time series models (ARIMA, GARCH) may struggle with non-linearities and exogenous factors. Econometric structural models rely on strong assumptions. While traditional machine learning (SVM, Random Forests) and standard deep learning models (LSTMs, CNNs) can capture complex patterns, they often function as ”black boxes,” limiting actionable insights [45]. There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
Given these limitations, this research addresses the question: How can we develop a deep learning framework that achieves high predictive accuracy for commodity futures prices across multiple horizons and simultaneously extracts a sparse, interpretable set of latent factors representing key market drivers? We propose a Regularized Sparse Autoencoder (RSAE) framework that utilizes an autoencoder backbone [96] with a time-lagged predictive objective and L1 sparsity regularization [94] on its latent code. This design, inspired by energy-based models [59] and sparse coding [72], forces the model to explain market dynamics through a parsimonious set of learned factors.
Our contribution lies in the specific design and application of this RSAE framework to multi-horizon commodity futures forecasting, aiming to enhance both accuracy and transparency. The remainder of this paper details the related work, the RSAE methodology, data, experimental setup, results, and discussion.
2 Review of Related Work in Commodity Forecasting
2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods—from agricultural produce to energy resources and industrial metals—affect many aspects of economic activities and individual well-being. For instance, volatile prices can significantly influence agricultural productivity and farmer incomes [35], while also presenting unpredictable expenses for consumers. Countries heavily dependent on commodity trade find their macroeconomic stability closely linked to commodity price fluctuations [18]. Additionally, amid rising concerns about global food security, effective prediction of staple commodity prices becomes vital to managing risks and preventing shortages [31].
Reliable commodity price forecasts are essential for various stakeholders. Producers depend on price predictions to make informed planting and harvesting decisions. Importers and exporters utilize forecasts to negotiate trade agreements and manage price risks effectively. Traders and investors in financial markets rely heavily on accurate price forecasts for speculative investments and portfolio management, viewing commodities as valuable assets that can diversify risk and potentially offer stability. Governments and policymakers require precise forecasts for macroeconomic planning, managing essential goods’ supply-demand dynamics, making informed decisions about agricultural support programs, and developing sound economic policies. This broad reliance highlights the systemic importance of commodity price fluctuations; inaccuracies in forecasting can have widespread economic consequences, impacting everything from consumer purchasing power to national economic performance and international trade dynamics.
Despite its significance, commodity price forecasting remains notably challenging. Commodity markets exhibit inherent complexity and volatility, characterized by nonlinear relationships and nonstationary trends [42]. Prices respond dynamically to a broad range of interrelated factors including fundamental supply-demand variables (production levels, crop yields, and inventory), macroeconomic indicators (inflation rates, industrial output, interest rates), financial market influences (currency exchange rates, investor behavior, volatility measures), geopolitical incidents, trade disruptions, and environmental variables like extreme weather events and climate change, especially affecting agricultural commodities. Furthermore, these influencing factors constantly evolve. Increased financial market involvement in commodities, frequent global crises (economic downturns, pandemics), erratic climate patterns, and shifting geopolitical conditions all add further complexity and unpredictability. The growing intricacy of these factors intensifies volatility, rendering forecasting increasingly challenging and demanding adaptive models capable of dynamically integrating new information and capturing evolving relationships. Consequently, traditional forecasting methods often fall short, necessitating exploration into more advanced modeling approaches.
2.2 Traditional Econometric Approaches: Foundations and Limitations
The quantitative forecasting of commodity prices has historically been dominated by traditional econometric and time series models. These methods provided the foundational tools for analyzing price dynamics and generating predictions based on historical data. Prominent among these are linear regression models, Autoregressive Integrated Moving Average (ARIMA) models, Vector Autoregression (VAR) models, Vector Error Correction Models (VECM) for cointegrated series, and Random Walk models. ARIMA models, in particular, gained widespread popularity over several decades, valued for their well-defined statistical properties and the systematic identification procedure proposed by Box and Jenkins [13]. Studies have demonstrated ARIMA’s utility in specific contexts, such as forecasting lead and zinc prices or certain agricultural commodity prices under specific conditions [68].
Recognizing the pervasive issue of time-varying volatility (i.e., periods of high volatility followed by periods of relative calm, known as volatility clustering) in financial and commodity markets, Generalized Autoregressive Conditional Heteroskedasticity (GARCH) models were developed [12, 29]. GARCH and its numerous extensions specifically aim to model and forecast the conditional variance of price returns. This focus on volatility acknowledges that risk assessment and management are critical aspects of commodity market analysis.
Despite their established role, traditional econometric models face significant limitations when applied to the complexities of modern commodity markets, primarily their reliance on restrictive assumptions like linearity and stationarity [90]. Commodity price series frequently exhibit strong nonlinearity and structural breaks [76], violating these core assumptions. While some studies report adequate performance for traditional models in specific scenarios [10], a growing body of evidence highlights their inadequacy for many commodity prices. The recognized shortcomings of these models served as a catalyst for exploring more flexible modeling paradigms [77, 56].
2.3 Rise of Machine Learning in Commodity Forecasting
In response to the limitations of traditional models, the field has seen a significant shift towards machine learning (ML) and artificial intelligence (AI) techniques [19, 85]. ML models can model complex, nonlinear relationships directly from data without restrictive assumptions, often leading to improved forecasting accuracy compared to traditional benchmarks [69]. ML also facilitates analysis of diverse datasets, including unstructured data via NLP [99].
This reflects a potential paradigm shift from theory-driven econometrics to data-driven ML [16]. However, ML models, especially deep learning, often require large datasets [37], which can be a constraint in some commodity markets.
2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied, as summarized in Table 1.
Table 1: Overview of Machine Learning Models in Commodity Forecasting.
Model Type Core Principle Strengths in Commodity Forecasting Weaknesses / Challenges
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, can be computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Can be prone to overfitting if not properly tuned, less interpretable than single trees.
Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons. Universal approximators, capture non-linear patterns. Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
Long Short-Term Memory (LSTM) Type of RNN with gating mechanisms to control information flow and capture long-range dependencies. Excels at modeling sequential data and long-term dependencies. Computationally expensive, requires significant data, complex architecture.
Gated Recurrent Unit (GRU) Simplified version of LSTM with fewer gates. Captures long-term dependencies like LSTM, potentially more computationally efficient. Performance relative to LSTM is context-dependent.
Transformer Uses self-attention mechanisms to weigh input sequence importance, parallelizable. Excellent at capturing long-range dependencies. Computationally very expensive, requires large datasets.
Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
Support Vector Regression (SVR) [27] and tree-based ensembles like Random Forests (RF) [15] and XGBoost [20] are foundational. SVR handles nonlinearities well [93, 50, 74, 73]. Tree-based methods are robust and can model interactions [92].
Deep learning, starting with ANNs/MLPs [84, 102], evolved to Recurrent Neural Networks (RNNs) [28] for sequential data. LSTMs [46] specifically address RNNs’ vanishing gradient problem [8], making them effective for long-term dependencies in commodity forecasting [30, 103, 88]. Architectural variants like bidirectional LSTMs [86] and CNN-LSTM hybrids [53, 65] exist. GRUs [21] offer a simpler alternative, with performance varying by context [100, 22]. CNNs [58, 7] also find use, often in hybrids. The choice often involves trade-offs between complexity, data needs, and interpretability [44].
Transformer models [95], with their self-attention mechanism, represent a newer frontier, excelling at long-range dependencies [104, 62, 61, 64]. They are often combined with decomposition techniques in hybrid models [97]. The development of very large models (LLMs) also influences this space [17, 98].
Hybrid models combining different techniques are common [101, 52]. A prevalent strategy is ”decomposition and ensemble” [57], using methods like Wavelet Transforms [24], VMD [26], EMD [47], SSA [36], or STL [23] to simplify the series before forecasting each component [6, 40, 43, 60]. Explainable AI (XAI) is also being integrated [75].
Performance evaluation uses metrics like RMSE, MAE, and directional accuracy [48], sometimes with statistical tests like Diebold-Mariano [25]. ML/DL models often outperform traditional ones, but rankings vary by context [91, 49].
3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
This section details the Regularized Sparse Autoencoder (RSAE) framework, an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from energy-based models [59] and sparse coding [72], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
3.1 Problem Formulation
We consider a set of $d$ market features (e.g., macroeconomic, supply/demand, technical indicators) at time $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input consists of a look-back window of $W$ historical observations:
$X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$
Our objective is to predict the commodity futures price $p_{t}$ at $N$ distinct future horizons, where $\tau_{1}<\tau_{2}<\ldots<\tau_{N}$. The corresponding target vector is:
$Y=Y_{t}=(p_{t+\tau_{1}},p_{t+\tau_{2}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$
The RSAE model aims to learn a mapping from $X$ to $Y$ via an intermediate sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$, where $m$ is the dimensionality of the latent space and $m\ll W\times d$. This latent variable $\mathbf{z}$ is intended to capture the essential underlying market state relevant for forecasting future price evolution.
3.2 Model Architecture
The RSAE architecture comprises three distinct neural network modules interacting within an energy-based learning paradigm. (A conceptual diagram of the architecture is presented in Figure 1.)
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework, illustrating the interaction between the History Predictor (Pred), Decoder (Dec), and Encoder (Enc) modules, along with the inference process for the sparse latent variable $\mathbf{z}$.
  1. 1.
    History Predictor (Pred): This module, parameterized by $\theta_{\text{Pred}}$, processes the historical input window $X$ to generate a condensed summary or context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$. Given the sequential nature of $X$, this module typically utilizes recurrent layers (e.g., LSTMs [46], GRUs [21]) or temporal convolutional networks (1D CNNs) [7].
    $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$
  2. 2.
    Decoder (Dec): This module, parameterized by $\theta_{\text{Dec}}$, takes both the history context $\bm{h}$ and a specific latent variable $\bm{z}\in\mathbb{R}^{m}$ as input to predict the multi-horizon future price vector $\tilde{Y}\in\mathbb{R}^{N}$.
    $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$
    The latent variable $\bm{z}$ acts as a representation of the underlying factors modulating the future trajectory predicted from the historical context $\bm{h}$.
  3. 3.
    Encoder (Enc): This module, parameterized by $\theta_{\text{Enc}}$, performs amortized inference [33, 55]. Its role is to learn to directly predict the optimal latent variable $\bm{z}^{*}$ (derived via the inference process detailed below) given the actual future outcome $Y$ and the history context $\bm{h}$.
    $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
    where $\hat{\bm{z}}$ is the Encoder’s estimate of $\bm{z}^{*}$.
3.3 Latent Variable Framework: Energy Function and Inference
Central to the RSAE is the determination of the latent variable $\bm{z}$. We define an energy function $E(Y,\bm{h},\bm{z})$ that quantifies the compatibility between a latent code $\bm{z}$, the historical context $\bm{h}$, and the future outcome $Y$. Low energy corresponds to a sparse latent code $\bm{z}$ that enables the Decoder to accurately predict $Y$ from $\bm{h}$ [59].
Energy Function
$E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$
where:
  • $C(Y,\tilde{Y})$ is the prediction cost, measuring the error between the true future $Y$ and the prediction $\tilde{Y}=\text{Dec}(\bm{z},\bm{h})$. We typically use Mean Squared Error (MSE) [11]:
    $C(Y,\tilde{Y})=\frac{1}{N}\left\|Y-\tilde{Y}\right\|_{2}^{2}$
  • $R(\bm{z})$ is the L1-norm sparsity regularizer [94]:
    $R(\bm{z})=\|\bm{z}\|_{1}=\sum_{k=1}^{m}|z_{k}|$
  • $\lambda>0$ is the hyperparameter controlling the strength of the sparsity penalty.
Optimal Latent Variable ($\bm{z}^{*}$)
For a given input pair $(X,Y)$ yielding $(\bm{h},Y)$, the optimal latent variable $\bm{z}^{*}$ is the value of $\bm{z}$ that minimizes the energy function:
$\bm{z}^{*}=\arg\min_{\bm{z}\in\mathbb{R}^{m}}E(Y,\bm{h},\bm{z})$
This minimization problem is typically non-convex if the decoder is non-linear.
Inference via Optimization
Finding $\bm{z}^{*}$ involves solving this optimization problem. We employ an iterative approach using gradient descent (or a subgradient method due to the L1 term) for a fixed number of steps $K$. This constitutes the explicit inference stage within the training loop:
  1. 1.
    Initialize $\bm{z}^{(0)}$ (e.g., a zero vector or from the Encoder).
  2. 2.
    For $k=0$ to $K-1$:
    • Compute the (sub)gradient of the energy $E$ with respect to $\bm{z}$ at $\bm{z}^{(k)}$, holding $\bm{h}$ and $\theta_{\text{Dec}}$ constant [14]:
      $\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})=\frac{\partial C}{\partial\text{Dec}}\cdot\frac{\partial\text{Dec}(\bm{z}^{(k)},\bm{h})}{\partial\bm{z}}+\lambda\,\text{sign}(\bm{z}^{(k)})$
    • Update $\bm{z}$ using a step size $\alpha_{\text{inf}}$:
      $\bm{z}^{(k+1)}\leftarrow\bm{z}^{(k)}-\alpha_{\text{inf}}\nabla_{\bm{z}}E(\cdot,\cdot,\bm{z}^{(k)})$
  3. 3.
    The result $\bm{z}^{(K)}$ serves as the inferred optimal latent variable for this sample: $\bm{z}^{*}\approx\bm{z}^{(K)}$.
This inference process finds a sparse latent code $\bm{z}^{*}$ under the current model parameters that aims to best explain the observed future $Y$ given the history $\bm{h}$.
3.4 Amortized Inference via Encoder (Enc)
The $K$-step inference process, while crucial for finding an optimal latent variable $\bm{z}^{*}$, introduces computational overhead during training and especially at test time if performed for every sample. To address this, the Encoder module $\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$ provides amortized inference by learning to directly map the observed future $Y$ and historical context $\bm{h}$ to an estimate $\hat{\bm{z}}$ of $\bm{z}^{*}$ [33].
$\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
This learned mapping enables faster inference once the model is trained. During training, the explicit $K$-step inference process is used to generate the target $\bm{z}^{*}$ for supervising the Encoder’s learning. At test time, for forecasting, the Encoder cannot use the future $Y$; instead, an alternative encoder $\text{Enc}_{\text{test}}(X;\theta_{\text{EncTest}})$ might be trained to predict $\bm{z}^{*}$ from $X$ only, or the $K$-step inference is run starting from $\bm{z}^{(0)}=0$. This paper uses the latter approach for generating $\bm{z}^{*}$ at test time.
3.5 Training Objective and Parameter Updates
The training process aims to minimize errors in both prediction (using the inferred latent variable $\bm{z}^{*}$) and in the amortized inference of $\bm{z}^{*}$ by the Encoder (if an Encoder for $\bm{z}^{*}$ from $X$ and $Y$ is used for training efficiency). The parameters $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{Enc}}$ are updated based on two primary loss components, calculated after obtaining $\bm{h}=\text{Pred}(X)$, inferring $\bm{z}^{*}$ via $K$-step optimization, and, if applicable, computing $\hat{\bm{z}}=\text{Enc}(Y,\bm{h})$:
  1. 1.
    Prediction Loss ($L_{\text{Pred}}^{*}$): Measures the error of the Decoder when using the inferred optimal latent variable $\bm{z}^{*}$.
    $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$
  2. 2.
    Inference Loss ($L_{\text{Inf}}$) (Optional, for training the Encoder Enc): Measures the discrepancy between the optimal latent variable $\bm{z}^{*}$ (from K-step inference) and the training-time Encoder’s prediction $\hat{\bm{z}}$.
    $L_{\text{Inf}}=D(\bm{z}^{*},\hat{\bm{z}})=\left\|\bm{z}^{*}-\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})\right\|_{2}^{2}$
The gradients for updating the network parameters are derived from these losses:
  • $\theta_{\text{Dec}}$ is updated based on $\nabla_{\theta_{\text{Dec}}}L_{\text{Pred}}^{*}$.
  • $\theta_{\text{Enc}}$ (if used) is updated based on $\nabla_{\theta_{\text{Enc}}}L_{\text{Inf}}$.
  • $\theta_{\text{Pred}}$ is updated based on gradients flowing back from $L_{\text{Pred}}^{*}$ (via $\bm{h}$) and potentially $L_{\text{Inf}}$ (via $\bm{h}$, if Enc is used). The dependency of $\bm{z}^{*}$ on $\bm{h}$ and $\theta_{\text{Dec}}$ can be handled by treating $\bm{z}^{*}$ as fixed post-inference for simplicity, or through more complex methods like implicit differentiation [79] or unrolling the inference loop [39]. In this work, $\bm{z}^{*}$ is treated as fixed after its $K$-step inference for updating $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$.
The sparsity penalty $\lambda\|\bm{z}\|_{1}$ directly shapes the target $\bm{z}^{*}$ during its $K$-step inference, thereby influencing all subsequent parameter updates.
3.6 Overall Training Algorithm
The training follows a mini-batch stochastic gradient descent procedure:
  1. 1.
    Initialization: Set up data loaders, initialize network parameters $(\theta_{\text{Pred}},\theta_{\text{Dec}})$, and choose hyperparameters ($W$, $m$, $N$, $\tau_{i}$, $\lambda$, $K$, $\alpha_{\text{inf}}$, batch size $B$, optimizer learning rate $\eta$). The Encoder $\theta_{\text{Enc}}$ is optional for the core prediction task if inference is always done via K-step optimization.
  2. 2.
    Epoch Loop: For each mini-batch $\{(X^{(j)},Y^{(j)})\}_{j=1}^{B}$:
    1. (a)
      Compute history encoding: $\bm{h}^{(j)}=\text{Pred}(X^{(j)};\theta_{\text{Pred}})$
    2. (b)
      Infer optimal latent variable via $K$-step gradient descent:
      $\bm{z}^{*(j)}=\arg\min_{\bm{z}}E(Y^{(j)},\bm{h}^{(j)},\bm{z})$
      (using current $\theta_{\text{Dec}}$)
    3. (c)
      Calculate per-sample prediction loss: $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)};\theta_{\text{Dec}}))$
    4. (d)
      Compute average batch prediction loss: $L_{\text{Pred}}^{*}=\frac{1}{B}\sum_{j=1}^{B}L_{\text{Pred}}^{*(j)}$
    5. (e)
      Compute gradients of $L_{\text{Pred}}^{*}$ with respect to $\theta_{\text{Pred}},\theta_{\text{Dec}}$.
    6. (f)
      Update parameters $\theta_{\text{Pred}},\theta_{\text{Dec}}$ using the chosen optimizer (e.g., Adam [54]).
  3. 3.
    Validation & Tuning: Periodically evaluate $L_{\text{Pred}}^{*}$ (using $\bm{z}^{*}$ inferred via $K$-step optimization at validation time) and sparsity metrics (e.g., $\|\bm{z}^{*}\|_{0}$) on a validation set. Use results to guide hyperparameter tuning (especially $\lambda$, $K$, $m$) and for early stopping [78].
  4. 4.
    Testing & Analysis: Evaluate the final trained model on a held-out test set using $L_{\text{Pred}}^{*}$ as the primary performance metric (with $\bm{z}^{*}$ inferred via $K$-step optimization). Analyze the inferred latent variables $\bm{z}^{*}$ from the test set—examining their sparsity, correlations with input features, and temporal evolution—to assess interpretability.
3.7 Interpretability via Sparsity
The explicit use of the L1 norm $R(\bm{z})=\|\bm{z}\|_{1}$ in the energy function plays a critical role in achieving interpretability [94]. By encouraging components of the optimal latent variable $\bm{z}^{*}$ to be exactly zero, the model is forced to represent the underlying market dynamics responsible for predicted price changes using a minimal set of active latent factors. Analyzing which factors are active can provide insights into the model’s reasoning. The hyperparameter $\lambda$ directly controls this sparsity.
4 Data Description and Pre-processing
4.1 Data Sources
Data was aggregated from multiple publicly accessible or commonly subscribed financial data provider sources:
  • Financial Information Providers: Services like Refinitiv Eikon or Bloomberg (via institutional access), and publicly available APIs from exchanges (e.g., CME Group, ICE) or financial data aggregators (e.g., Quandl - now part of Nasdaq Data Link, Yahoo Finance) — for daily market prices, futures data, and economic indicators.
  • Commodity Exchanges: Direct data feeds or historical data services from London Metal Exchange (LME), CME Group (NYMEX), and Intercontinental Exchange (ICE).
  • Government Agencies: U.S. Energy Information Administration (EIA), Bureau of Labor Statistics (BLS), Federal Reserve Economic Data (FRED), and National Bureau of Statistics of China.
  • International Organizations: International Monetary Fund (IMF), World Bank, Organisation for Economic Co-operation and Development (OECD).
  • Specialized Data Providers: Baker Hughes (rig counts via public releases), Commodity Futures Trading Commission (CFTC) (Commitment of Traders reports), and NOAA/NWS or equivalent national/international meteorological services (for historical weather data).
While the primary price data for model training is daily, several fundamental indicators are incorporated at lower frequencies. Table 2 illustrates sources for some representative monthly variables.
Table 2: Illustrative sources for some monthly input variables used in this research.
Variables Period Unit Data source Data availability
Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. Energy Information Administration
(https://www.eia.gov/dnav/pet/hist/RBRTED.htm)
Freely available
Copper Price (benchmark) Monthly \$/ton World Bank Commodity Price Data (Pink Sheet)
(https://www.worldbank.org/en/research/commodity-markets)
Freely available
Copper Supply, Demand Indicators Monthly Various U.S. Geological Survey (USGS) Mineral Commodity Summaries
(https://www.usgs.gov/centers/national-minerals-information-center/mineral-commodity-summaries)
Freely available
Crude Oil Supply and Demand Indicators Monthly Various U.S. Energy Information Administration (EIA) Short-Term Energy Outlook (STEO) Freely available
4.2 Feature Selection
A wide array of features were selected, categorized as:
4.2.1 Target Variable (Daily Futures Prices)
  • Copper: LME Official Cash Settlement Price (or nearby futures contract).
  • Crude Oil: WTI front-month futures price (NYMEX); Brent Crude front-month futures (ICE).
4.2.2 Lagged Target Variables & Price Dynamics (Daily)
Historical prices, log returns, realized volatility [4], futures curve spreads [38].
4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest rates (central bank rates, sovereign yields like US 10-Year Treasury), currency rates (DXY, EUR/USD, USD/CNY).
4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories (LME, COMEX, SHFE), publicly reported global/major producer output, China’s consumption/imports, scrap indicators (where publicly accessible). Crude Oil: EIA weekly inventories, US production, publicly available OPEC+ figures, Baker Hughes rig counts, IEA/OPEC demand estimates.
4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, major equity indices (S&P 500, FTSE 100, Shanghai Composite).
4.2.6 Technical Indicators (Calculated from Daily Prices)
Moving Averages (SMA, EMA), Oscillators (RSI, MACD), ADX [70].
4.2.7 Other (Exploratory)
Inter-commodity spreads, relevant weather indices (e.g., ENSO, hurricane impact assessments if data accessible).
4.3 Data Frequency and Time Period
Primary modeling frequency is daily. Lower frequency data (weekly, monthly, quarterly) are forward-filled, strictly aligning with information release dates to prevent lookahead bias [34]. Particular care is taken with lagged official statistics (e.g., STEO) to ensure only historically available information is used at each time step $t$. The dataset spans January 1, 2005 – December 31, 2023.
4.4 Data Pre-processing
Standard steps are applied [37]:
  1. 1.
    Missing Data: Features with $\geq$40missing values removed; remaining gaps primarily forward-filled.
  2. 2.
    Feature Transformation: Log returns or percentage changes for non-stationary series. Log transformation for skewed positive features.
  3. 3.
    Normalization/Standardization: Z-score normalization, with parameters derived only from the training set.
  4. 4.
    Outlier Treatment: Applied cautiously (e.g., winsorization at 1st/99th percentiles) to preserve genuine volatility.
4.5 Final Dataset Construction
Input sequences $X_{t}\in\mathbb{R}^{W\times d}$ (look-back $W=60$ days) and targets $Y_{t}\in\mathbb{R}^{N}$ (log prices at $N=3$ horizons: $\tau_{1}=1,\tau_{2}=5,\tau_{3}=22$ days) are formed. Chronological splits: Training (2005–2018), Validation (2019–2020), Test (2021–2023) [9].
5 Experimental Setup and Results
5.1 Experimental Setup
RSAE Configuration
Implemented using Python and JAX.
  • Input/Output: $W=60$ days, $d\approx 50-70$ features. $Y_{t}=(\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22}))^{\top}$.
  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [89]). Dec & Enc (training only): MLPs (hidden layers 64, 32 units; ReLU). Latent dimension $m=16$.
  • Inference: $K=10$ gradient steps, $\alpha_{\text{inf}}=0.01$. At test time, inference starts from $\bm{z}^{(0)}=0$.
  • Sparsity: $\lambda$ tuned via grid search ($\{10^{-5},\dots,10^{-2}\}$) on validation set; $\lambda=5\times 10^{-4}$ (Copper), $\lambda=10^{-4}$ (Crude Oil).
  • Training: Adam optimizer [54], $\eta=10^{-4}$, batch $B=64$, max 100 epochs, early stopping (patience 10 on $L_{\text{Pred}}^{*}$) [78].
Baseline Models
Persistence, ARIMA (orders via ‘pmdarima‘ [2, 87]), XGBoost [20] (tuned via CV, separate model per horizon), Standard LSTM (similar to RSAE’s Pred, direct output).
Evaluation Metrics
RMSE, MAE, Directional Accuracy (DA) on log prices [48]. Average $L_{0}$ norm for RSAE sparsity.
5.2 Predictive Accuracy Results
Performance on the test set (2021–2023) is in Table 3 and Table 4.
Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
[…]
RSAE 1-day 0.0125 0.0093 58.2
5-day 0.0283 0.0210 55.1
22-day 0.0591 0.0452 53.5
Persistence 1-day $0.0152$ $0.0115$ $50.0$
5-day $0.0340$ $0.0258$ $50.0$
22-day $0.0705$ $0.0540$ $50.0$
ARIMA 1-day $0.0149$ $0.0112$ $51.5$
5-day $0.0335$ $0.0251$ $50.8$
22-day $0.0690$ $0.0525$ $50.3$
XGBoost 1-day $0.0135$ $0.0101$ $56.5$
5-day $0.0305$ $0.0228$ $54.0$
22-day $0.0630$ $0.0481$ $52.1$
LSTM 1-day $0.0130$ $0.0098$ $57.3$
5-day $0.0295$ $0.0220$ $54.5$
22-day $0.0615$ $0.0469$ $52.8$
Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
[…]
RSAE 1-day 0.0185 0.0139 59.1
5-day 0.0410 0.0315 56.3
22-day 0.0835 0.0650 54.2
Persistence 1-day $0.0225$ $0.0170$ $50.0$
5-day $0.0500$ $0.0385$ $50.0$
22-day $0.1020$ $0.0805$ $50.0$
ARIMA 1-day $0.0220$ $0.0165$ $51.2$
5-day $0.0490$ $0.0375$ $50.6$
22-day $0.1000$ $0.0780$ $50.2$
XGBoost 1-day $0.0198$ $0.0150$ $57.5$
5-day $0.0445$ $0.0340$ $55.1$
22-day $0.0890$ $0.0695$ $52.9$
LSTM 1-day $0.0192$ $0.0145$ $58.3$
5-day $0.0430$ $0.0328$ $55.7$
22-day $0.0865$ $0.0675$ $53.5$
Observations from Accuracy Results: The RSAE framework generally achieves competitive or superior predictive performance compared to the baselines.
  • Error Metrics (RMSE, MAE): RSAE consistently yields the lowest RMSE/MAE values across horizons for both commodities, suggesting higher accuracy in predicting price change magnitudes. Improvements are more noticeable at shorter horizons.
  • Directional Accuracy (DA): RSAE also shows strong DA, often outperforming baselines. ML-based models significantly outperform Persistence and ARIMA in DA.
  • Performance Degradation: As expected, accuracy decreases for all models as the forecast horizon extends.
  • Sparsity of RSAE: The average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil) out of $m=16$, indicating effective sparsity from L1 regularization.
These results suggest RSAE’s architecture can improve forecasting accuracy while enabling interpretability.
5.3 Interpretability Analysis of Latent Factors
Analysis of $\bm{z}^{*}$ on the test set provided insights.
Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active, varying over time, suggesting adaptive representation.
Correlation with Input Features and Qualitative Economic Interpretation
Correlations between active $\bm{z}^{*}_{k}$ and key input features informed qualitative interpretations. Copper (Example Interpretations):
  • $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory): Positively correlated with Global PMI/Chinese industrial production, negatively with LME inventories. Active during economic acceleration.
  • $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD): Positively correlated with DXY/US Treasury yields. Stronger dollar/rates often showed this factor active with downward price pressure.
  • $z^{*}_{Cu,3}$ (Supply Shock Indicator): Activated during reported supply disruptions (e.g., mine strikes), often coinciding with short-term rallies.
  • $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): Correlated with VIX and COT speculative positioning.
Crude Oil (Example Interpretations):
  • $z^{*}_{Oil,1}$ (US Demand & Inventory): Correlated with EIA US inventories (negatively with builds) and mobility/gasoline demand.
  • $z^{*}_{Oil,2}$ (Geopolitical Risk & OPEC+ Policy): Spiked around OPEC+ announcements and geopolitical events.
  • $z^{*}_{Oil,3}$ (Market Risk Appetite): Correlated with credit spreads, equity shifts, and OVX.
  • $z^{*}_{Oil,4}$ (Global ex-US Demand): Correlated with global growth forecasts and non-US industrial activity.
Case Study Examples
  • Russia-Ukraine War (Early 2022): For Crude Oil, $z^{*}_{Oil,2}$ increased dramatically, aligning with the oil price rally and geopolitical risks.
  • Recession Fears/China Lockdowns (Mid-Late 2022): For Copper, $z^{*}_{Cu,1}$ turned significantly negative during high recession concerns and China’s lockdowns, coinciding with copper price weakness.
These examples illustrate how RSAE’s sparse factors can align with market narratives.
5.4 Discussion of Results
The RSAE framework achieves competitive predictive accuracy while inducing a sparse latent representation. The interpretability analysis, though qualitative, suggests learned factors capture meaningful economic influences (e.g., ”Global Demand,” ”Geopolitical Risk”), a key advantage over black-box models. RSAE navigates the accuracy-interpretability trade-off by incorporating L1 sparsity in its energy-based formulation. These findings position RSAE as a promising approach for interpretable commodity futures forecasting.
6 Discussion and Conclusion
This study introduced and evaluated a Regularized Sparse Autoencoder (RSAE) framework for multi-horizon commodity futures forecasting and interpretable factor discovery. The experimental results suggest its potential.
6.1 Implications of Findings
  1. 1.
    Bridging Accuracy-Interpretability: RSAE shows a path to competitive predictive performance with transparency via sparse latent factors, vital in finance [83].
  2. 2.
    Enhanced Decision Support: Interpretable factors (e.g., ”Global Demand”) can aid hedging, inventory, and investment decisions.
  3. 3.
    Automated Factor Discovery: RSAE offers data-driven discovery of latent drivers and potential regime shift indicators [5].
  4. 4.
    Potentially Improved Risk Management: Validated sparse factors could enhance factor-based VaR or stress testing.
  5. 5.
    Broader Applicability: The architecture may apply to other financial/economic forecasting.
6.2 Limitations of the Study
  1. 1.
    Interpretability Validation: Factor interpretation is qualitative and post-hoc; rigorous quantitative validation (event studies, causality tests) is needed [63].
  2. 2.
    Input Feature Dependence: Model learning is constrained by input feature quality and comprehensiveness.
  3. 3.
    Model Complexity/Cost: RSAE’s multi-component structure and inference loop increase training time and tuning effort.
  4. 4.
    Latent Variable Inference: Fixed $K$-step inference is an approximation; convergence and optimality are not guaranteed across all scenarios. Adaptive or convergence-based criteria for $\bm{z}^{*}$ inference are important future considerations.
  5. 5.
    Structural Breaks/Non-Stationarity: Adaptability to unforeseen structural breaks not in training data is a challenge [82].
  6. 6.
    Sparsity Level Selection: Choosing $\lambda$ involves a trade-off between interpretability and accuracy.
  7. 7.
    No Intrinsic Uncertainty Quantification: The model provides point forecasts, lacking probabilistic output [51].
  8. 8.
    Experimental Design: A single chronological train/validation/test split was used; more robust time-series cross-validation (e.g., rolling-origin or blocked CV) would provide more reliable out-of-sample performance estimates. Baselines received standard tuning (e.g., ‘$auto\_arima$‘, single CV for XGBoost), while RSAE had a specific grid search for $\lambda$; a fully balanced hyperparameter tuning budget across all models would strengthen comparisons. The directional accuracy metric employed is basic; more advanced metrics like the Pesaran-Timmermann test or Matthews Correlation Coefficient should be considered. The absence of statistical significance tests (e.g., Diebold-Mariano) for performance differences is a limitation.
6.3 Avenues for Future Research
  1. 1.
    Advanced Interpretability/Validation: Integrate SHAP [66] or LIME [81]; use NLP on news to label/validate factors.
  2. 2.
    Improved Inference: Explore variational inference [55, 80] or adaptive optimization for $\bm{z}^{*}$.
  3. 3.
    Dynamic Sparsity Control: Allow $\lambda$ to be learned or vary with market context.
  4. 4.
    Regime-Switching Integration: Combine RSAE with Markov Switching Models [41].
  5. 5.
    Uncertainty Quantification: Extend to probabilistic forecasts (Bayesian NNs [71, 67], MC dropout [32]).
  6. 6.
    Hierarchical/Cross-Asset Modeling: Model related assets jointly to find shared drivers.
  7. 7.
    End-to-End Differentiable Optimization: Replace K-step inference with differentiable optimization layers [3, 39].
  8. 8.
    High-Frequency/Online Learning: Adapt for intraday data and online updates.
  9. 9.
    Reproducibility: Release code, preprocessed data, and random seeds to facilitate replication and extension.
  10. 10.
    Advanced Benchmarking: Compare against latest Transformer architectures (e.g., PatchTST) and apply rigorous statistical tests for performance differences (e.g., Diebold-Mariano tests, Model Confidence Sets). Evaluate with alternative directional accuracy metrics (e.g., Pesaran-Timmermann test).
6.4 Conclusion
This paper introduced the RSAE framework, a deep learning approach combining established techniques for multi-horizon commodity futures forecasting and interpretable latent factor discovery. By integrating time-lagged prediction with L1-regularized sparse inference, RSAE aims to balance predictive accuracy with model transparency. Empirical evaluation on Copper and Crude Oil futures data showed RSAE’s potential for competitive accuracy and for yielding sparse latent factors that correlate qualitatively with known market drivers. This offers an advantage over opaque black-box models. While acknowledging current limitations, including the qualitative nature of factor interpretation and the need for more extensive experimental validation (e.g., robust CV, statistical significance testing, broader hyperparameter tuning for baselines), the RSAE framework represents a promising direction for developing more insightful forecasting systems in finance. Future work focusing on advanced inference, uncertainty quantification, and rigorous validation is key to realizing its full potential.