### v3

# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

Abstract
Commodity futures price volatility creates significant economic challenges, necessitating accurate multi-horizon forecasting. Predicting these prices is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity futures prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. A key L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand shifts, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings suggest the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a notable advantage over traditional black-box approaches.
1 Introduction
Commodity markets, and their associated futures contracts, form a critical part of the global economy, providing essential raw materials and hedging instruments. The prices of key commodity futures, such as those for industrial metals like Copper and energy benchmarks like Crude Oil, are not merely transactional figures; they are vital indicators of economic health, drivers of industrial costs, and significant factors influencing inflation, trade balances, and geopolitical dynamics. Copper futures, for instance, are closely watched for signals about global industrial demand. Crude Oil futures are pivotal for managing energy price risk and reflect complex geopolitical and supply/demand interactions.
The inherent volatility of these commodity futures prices presents substantial risks and opportunities for a wide spectrum of market participants. Producers and consumers use futures to hedge price uncertainty, while financial institutions and traders engage in speculative and arbitrage activities. Consequently, the ability to accurately forecast commodity futures prices across various time horizons is of paramount importance for effective risk mitigation, optimized resource allocation, and informed investment decisions.
The challenge lies in the sheer complexity of factors driving futures prices. These markets are influenced by a confluence of interconnected forces:
  1. 1.
    Macroeconomic Environment: Global GDP growth, industrial production, interest rates, currency exchange rates (particularly the US Dollar), and inflation expectations.
  2. 2.
    Underlying Physical Market Fundamentals: For Copper, this includes mine production, scrap availability, and inventory levels. For Crude Oil, factors include OPEC+ decisions, non-OPEC output, geopolitical events in producing regions, and inventory levels.
  3. 3.
    Demand Fundamentals: Linked to manufacturing activity, construction, transportation needs, and energy transition dynamics.
  4. 4.
    Market Sentiment and Financialization: Speculative positioning in futures and options markets (e.g., CFTC Commitment of Traders reports), algorithmic trading, and broader risk-on/risk-off sentiment.
  5. 5.
    Other factors: Including weather patterns impacting supply/demand and inter-commodity relationships.
This intricate web of high-dimensional, often collinear, and dynamically interacting drivers makes futures price forecasting an exceptionally challenging task, characterized by non-linear relationships and regime shifts. Existing approaches often face a trade-off between predictive accuracy and interpretability. Classical time series models (ARIMA, GARCH) may struggle with non-linearities and exogenous factors. Econometric structural models rely on strong assumptions. While traditional machine learning (SVM, Random Forests) and standard deep learning models (LSTMs, CNNs) can capture complex patterns, they often function as ”black boxes,” limiting actionable insights [47]. There is a need for models that not only forecast accurately but also provide an understanding of the underlying drivers.
Given these limitations, this research addresses the question: How can we develop a deep learning framework that achieves high predictive accuracy for commodity futures prices across multiple horizons and simultaneously extracts a sparse, interpretable set of latent factors representing key market drivers? We propose a Regularized Sparse Autoencoder (RSAE) framework that utilizes an autoencoder backbone [99] with a time-lagged predictive objective and L1 sparsity regularization [97] on its latent code. This design, inspired by energy-based models [61] and sparse coding [75], forces the model to explain market dynamics through a parsimonious set of learned factors.
Our contribution lies in the specific design and application of this RSAE framework to multi-horizon commodity futures forecasting, aiming to enhance both accuracy and transparency. The remainder of this paper details the related work, the RSAE methodology, data, experimental setup, results, and discussion.
2 Review of Related Work in Commodity Forecasting
2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods affect economic activities and individual well-being [36]. Macroeconomic stability of commodity-dependent countries is closely linked to price fluctuations [19], and predicting staple commodity prices is vital for food security [32]. Reliable forecasts are essential for producers, traders, investors, and policymakers.
Commodity price forecasting is notably challenging due to inherent market complexity, volatility, nonlinear relationships, and nonstationary trends [43]. Prices respond to numerous interrelated factors including supply/demand fundamentals, macroeconomic indicators, financial market influences, geopolitical events, and environmental variables. These factors evolve, with increased financialization, global crises, and climate change adding unpredictability, often rendering traditional methods inadequate.
2.2 Traditional Econometric Approaches: Foundations and Limitations
Quantitative commodity forecasting historically relied on models like ARIMA, VAR, and VECM. Box-Jenkins ARIMA models [14] gained popularity for their statistical properties [70]. GARCH models [13, 30] were developed to address time-varying volatility. However, these models struggle with the nonlinearities and structural breaks common in commodity prices [93, 80], often underperforming [11]. Their limitations spurred exploration of more flexible paradigms [81, 58].
2.3 Rise of Machine Learning in Commodity Forecasting
ML and AI techniques have become prominent alternatives [20, 88]. ML models inherently handle complex, nonlinear relationships without restrictive assumptions, often improving accuracy over traditional methods [71]. They can also integrate diverse data, including unstructured text via NLP [102]. This data-driven approach [17] contrasts with theory-driven econometrics but typically requires large datasets, a potential constraint in some markets [38].
2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied (Table 1).
Table 1: Overview of Machine Learning Models in Commodity Forecasting.
Model Type Core Principle Strengths in Commodity Forecasting Weaknesses / Challenges
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Prone to overfitting if not properly tuned, less interpretable than single trees.
Artificial Neural Network (ANN) / Multilayer Perceptron (MLP) Learns complex non-linear mappings through interconnected layers of neurons. Universal approximators, capture non-linear patterns. Prone to local optima, requires careful tuning, can overfit, may struggle with long-term dependencies.
Long Short-Term Memory (LSTM) Type of RNN with specialized gating mechanisms (input, forget, output gates) to control information flow and capture long-range dependencies in sequential data. Excels at modeling sequential data and long-term dependencies, mitigating vanishing gradient problem of simple RNNs. Computationally expensive, requires significant data, complex architecture demanding careful tuning.
Gated Recurrent Unit (GRU) Simplified LSTM version with fewer gates (reset and update gates), also for long-range dependencies. Captures long-term dependencies like LSTM, potentially more computationally efficient due to fewer parameters. Performance relative to LSTM is context-dependent, may not always match LSTM’s capabilities on complex tasks.
Transformer Uses self-attention mechanisms to weigh input sequence importance globally, allowing parallel processing of sequence elements. Excellent at capturing very long-range dependencies, often outperforming RNNs/LSTMs on such tasks; parallelizable. Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
Note: Strengths, weaknesses, and representative studies are derived from the provided research material.
Foundational models include SVR [28], noted for handling non-linearities [96, 52, 77, 76], and tree-based ensembles like RF [16] and XGBoost [21], valued for robustness and interaction modeling [95].
Deep learning (DL) advanced from ANNs/MLPs [87, 105] to RNNs [29]. LSTMs [48] and GRUs [22] became key for sequential data due to their ability to learn long-term dependencies by addressing the vanishing gradient issue [8]. They are widely used in commodity forecasting [31, 106, 91, 103, 23], with variants like bidirectional LSTMs [89] and CNN-LSTM hybrids [55, 67, 60, 7]. The choice involves balancing complexity, data needs, and interpretability [46].
Transformer models [98], leveraging self-attention, represent a significant advancement for long-range dependency capture [64, 63, 66]. Recent architectures like PatchTST [74] and FEDformer [107] highlight ongoing progress. Transformers are often integrated into hybrid models, frequently with decomposition methods [100]. Large Language Models (LLMs) are also influencing financial forecasting research [18, 101].
Hybrid models are prevalent, often combining traditional methods with ML, or using a ”decomposition and ensemble” strategy [104, 54, 59]. Decomposition techniques (e.g., Wavelets [25], VMD [27], EMD [49], SSA [37], STL [24]) simplify series for component-wise forecasting [6, 41, 45, 62]. Integration of XAI is also an emerging trend [78].
Performance is typically assessed using RMSE, MAE, and directional accuracy [50]. Statistical tests like Diebold-Mariano [26] are sometimes used for comparing model accuracy. While ML/DL models often show superior performance, rankings are context-dependent [94, 51].
3 Methodology: The Regularized Sparse Autoencoder (RSAE) Framework
This section details the Regularized Sparse Autoencoder (RSAE) framework, an approach specifically designed for the dual task of multi-horizon commodity futures price prediction and the discovery of interpretable latent market factors. The methodology draws inspiration from energy-based models [61] and sparse coding [75], employing an explicit inference step to determine a sparse latent representation that best explains future price movements.
3.1 Problem Formulation
We consider a set of $d$ market features at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input is a look-back window $X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$. Our objective is to predict the futures price $p_{t}$ at $N$ horizons $\tau_{1}<\ldots<\tau_{N}$, with target $Y=Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$. The RSAE learns a mapping from $X$ to $Y$ via a sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$ ($m\ll W\times d$).
3.2 Model Architecture
The RSAE comprises three neural network modules (Figure 1).
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework.
  1. 1.
    History Predictor (Pred): $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$, maps $X$ to a context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$, using LSTMs [48], GRUs [22], or 1D CNNs [7].
  2. 2.
    Decoder (Dec): $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$, predicts future prices $\tilde{Y}\in\mathbb{R}^{N}$ from $\bm{z}$ and $\bm{h}$.
  3. 3.
    Encoder (Enc) (for training): $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$, performs amortized inference of $\bm{z}^{*}$ from actual future $Y$ and $\bm{h}$ [34, 57]. A separate encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, is trained to predict $\bm{z}^{*}$ from $X$ alone for test-time forecasting. $\text{Enc}_{\text{XtoZ}}$ is trained with a mean-squared error objective to regress the K-step optimal z*; its parameters are not updated through the prediction loss.
3.3 Latent Variable Framework: Energy Function and Inference
We define an energy function $E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$ [61]. $C(Y,\tilde{Y})=\frac{1}{N}\|Y-\tilde{Y}\|_{2}^{2}$ is the MSE prediction cost [12]. $R(\bm{z})=\|\bm{z}\|_{1}$ is the L1-norm sparsity regularizer [97]. The optimal latent variable $\bm{z}^{*}=\arg\min_{\bm{z}}E(Y,\bm{h},\bm{z})$ is found via $K$-step (sub)gradient descent [15].
3.3.1 Selection and Convergence of Inference Steps (K)
The number of iterative steps, $K$, in the inference process for $\bm{z}^{*}$ is a hyperparameter balancing thoroughness of energy minimization against computational cost. To determine an appropriate value for $K$, preliminary experiments were conducted on the validation set. This involved tracking the average energy function value, $E(Y,\bm{h},\bm{z})$, over a range of potential iterations (e.g., $k=1$ to $k=15$) for multiple representative batches of validation samples.
The observed convergence behavior is depicted in Figure 2. The mean energy (averaged across validation batches) typically exhibited a sharp decrease during the initial iterations (approximately $k=1$ to $k=7$). Beyond this initial phase, the rate of energy reduction substantially diminished, with the curve beginning to plateau around $k=8$ to $k=10$. While formal convergence to a global minimum is not guaranteed for non-convex energy landscapes, the iterative process aims to find a suitable local minimum that provides a sparse and predictive latent code. Based on the empirical observation that iterations beyond $K=10$ provided only marginal decreases in the objective energy at an increased computational cost, $K=10$ was selected for all subsequent experiments in this study. A more extensive analysis of convergence properties under varying conditions remains an avenue for future work.
Figure 2: Mean objective energy function value $E(Y,\bm{h},\bm{z})$ versus the number of inference iterations ($k$), averaged over representative validation set batches. The plot demonstrates typical convergence behavior, with energy decreasing rapidly initially before plateauing. The selected $K=10$ (dashed red line) represents a point of diminishing returns for energy reduction.
3.4 Test-Time Inference
At test time, future $Y$ is unknown. We train an additional inference encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, supervised by the $\bm{z}^{*}$ values obtained during the training phase (from the $K$-step optimization using known $Y_{\text{train}}$). Thus, for a new test input $X_{\text{test}}$, the latent code is obtained directly: $\bm{z}_{\text{test}}=\text{Enc}_{\text{XtoZ}}(X_{\text{test}};\theta_{\text{XtoZ}})$. This amortized inference from $X$ alone is used for generating forecasts.
3.5 Training Objective and Parameter Updates
The primary training objective minimizes the prediction loss using the $K$-step inferred $\bm{z}^{*}$: $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$. Parameters $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$ are updated based on gradients from $L_{\text{Pred}}^{*}$. If the training-time Encoder $\text{Enc}(Y,\bm{h})$ is used, it is trained with an inference loss $L_{\text{Inf}}=\|\bm{z}^{*}-\text{Enc}(Y,\bm{h})\|_{2}^{2}$. The test-time encoder $\text{Enc}_{\text{XtoZ}}(X)$ is trained separately to predict $\bm{z}^{*}$ from $X$.
3.6 Overall Training Algorithm
  1. 1.
    Initialize $\theta_{\text{Pred}},\theta_{\text{Dec}},\theta_{\text{XtoZ}}$.
  2. 2.
    For each training batch $\{(X^{(j)},Y^{(j)})\}$:
    1. (a)
      $\bm{h}^{(j)}=\text{Pred}(X^{(j)})$.
    2. (b)
      Infer $\bm{z}^{*(j)}$ via $K$-step minimization of $E(Y^{(j)},\bm{h}^{(j)},\bm{z})$.
    3. (c)
      $L_{\text{Pred}}^{*(j)}=C(Y^{(j)},\text{Dec}(\bm{z}^{*(j)},\bm{h}^{(j)}))$.
    4. (d)
      Update $\theta_{\text{Pred}},\theta_{\text{Dec}}$ via gradients of $\text{mean}(L_{\text{Pred}}^{*})$ (e.g., Adam [56]).
    5. (e)
      Update $\theta_{\text{XtoZ}}$ to predict $\bm{z}^{*(j)}$ from $X^{(j)}$.
  3. 3.
    Validate using $L_{\text{Pred}}^{*}$ (with $\bm{z}_{\text{val}}=\text{Enc}_{\text{XtoZ}}(X_{\text{val}})$) for tuning and early stopping [82].
  4. 4.
    Test similarly. Analyze inferred $\bm{z}_{\text{test}}$ for interpretability.
3.7 Interpretability via Sparsity
The L1 norm encourages sparse $\bm{z}^{*}$, aiding interpretability [97].
4 Data Description and Pre-processing
4.1 Data Sources
Aggregated from public/common sources: Refinitiv Eikon, Bloomberg, exchange APIs (CME, ICE, LME), government agencies (EIA, BLS, FRED, NBS China), international organizations (IMF, World Bank, OECD), and specialized providers (Baker Hughes, CFTC, NOAA/NWS). Table 2 gives examples.
Table 2: Illustrative sources for some monthly input variables.
Variables Period Unit Data source Data availability
Crude Oil Price (benchmark, e.g., Brent) Monthly \$/barrel U.S. EIA (Pet. Hist.) Freely available
Copper Price (benchmark) Monthly \$/ton World Bank (Comm. Data) Freely available
Copper Supply, Demand Ind. Monthly Various USGS (Min. Sum.) Freely available
Crude Oil Supply, Demand Ind. Monthly Various U.S. EIA Short-Term Energy Outlook (STEO) Freely available
4.2 Feature Selection
4.2.1 Target Variable (Daily Futures Prices)
Copper (LME nearby futures), Crude Oil (WTI, Brent front-month).
4.2.2 Lagged Target Variables & Price Dynamics (Daily)
Historical prices, log returns, realized volatility [4], futures curve spreads [39].
4.2.3 Macroeconomic Indicators (Daily, Monthly, Quarterly)
Global/Major Economy PMIs, Industrial Production, inflation (CPI, PPI), interest/currency rates.
4.2.4 Supply/Demand Fundamentals (Varying Frequencies)
Copper: Exchange inventories, producer output, China consumption/imports, scrap indicators. Crude Oil: EIA weekly inventories, US/OPEC+ production, rig counts, demand estimates.
4.2.5 Market Sentiment & Financial Data (Daily, Weekly)
Futures volume/open interest, CFTC COT reports, VIX, OVX, S&P 500, FTSE 100, Shanghai Comp.
4.2.6 Technical Indicators (Calculated from Daily Prices)
SMAs, EMAs, RSI, MACD, ADX [72].
4.2.7 Other (Exploratory)
Inter-commodity spreads, weather indices (ENSO, hurricane impacts).
4.3 Data Frequency and Time Period
Primary modeling is daily. Lower frequency data forward-filled, with release lags respected for official stats (e.g., STEO) to prevent lookahead bias [35]. Dataset: Jan 1, 2005 – Dec 31, 2023.
4.4 Data Pre-processing
Standard steps [38]: Missing data handling (remove if $\geq$40%, else ffill), transformations for stationarity/skewness (log returns, log transform), Z-score normalization (params from training set only), cautious outlier treatment.
4.5 Final Dataset Construction
Inputs $X_{t}\in\mathbb{R}^{W\times d}$ ($W=60$), targets $Y_{t}\in\mathbb{R}^{N}$ ($N=3$ horizons: 1, 5, 22 days). Chronological splits: Train (2005–2018), Valid (2019–2020), Test (2021–2023) [9]. Robustness could be improved with rolling-origin CV [10].
5 Experimental Setup and Results
The empirical results presented herein are based on the models trained and evaluated according to the setup detailed below.
5.1 Experimental Setup
RSAE Configuration
Implemented using Python and JAX.
  • Input/Output: $W=60$ days, $d\approx 50-70$ features. $Y_{t}=(\log(p_{t+1}),\log(p_{t+5}),\log(p_{t+22}))^{\top}$.
  • Architecture: Pred: 2 stacked LSTMs (128 units each, dropout 0.2 [92]). Dec: MLP (hidden layers 64, 32 units; ReLU). A separate encoder $\text{Enc}_{\text{XtoZ}}(X)$ maps $X$ to $\bm{z}^{*}$ at test time. Latent dimension $m=16$.
  • Inference (Training): $K=10$ gradient steps for $\bm{z}^{*}$, $\alpha_{\text{inf}}=0.01$.
  • Sparsity: $\lambda$ tuned via grid search ($\{10^{-5},\dots,10^{-2}\}$) on validation set; $\lambda=5\times 10^{-4}$ (Copper), $\lambda=10^{-4}$ (Crude Oil).
  • Training: Adam optimizer [56], $\eta=10^{-4}$, batch $B=64$, max 100 epochs, early stopping (patience 10 on $L_{\text{Pred}}^{*}$) [82].
Baseline Models
Persistence, ARIMA [14] (orders via ‘pmdarima‘ [2, 90]), XGBoost [21] (tuned via CV), Standard LSTM. Baseline hyperparameter tuning was standard; a fully balanced budget vs. RSAE is a future work consideration.
Evaluation Metrics
Forecast performance was evaluated using standard metrics on log-transformed prices [50]: Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE). For directional performance, we report several metrics:
  • Standard Directional Accuracy (DA): The percentage of times the model correctly predicts the direction of price change from the current price $p_{t}$.
    $DA_{\text{std}}=\frac{1}{T_{\text{test}}}\sum_{t\in\text{Test}}\mathbb{I}\left[\text{sign}(\hat{p}_{t+\tau_{i}}-p_{t})=\text{sign}(p_{t+\tau_{i}}-p_{t})\right]$
    where $\text{sign}(0)$ is treated as 0. For this calculation, if both actual and predicted change are zero (within a small tolerance, $\epsilon_{\text{zero}}$), it is counted as a correct prediction.
[…]
  • Up-Move Hit Rate (UHR) & Down-Move Hit Rate (DHR): Calculated on the subset of days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). UHR is the proportion of actual up-moves correctly predicted as up-moves. DHR is the proportion of actual down-moves correctly predicted as down-moves.
  • Matthews Correlation Coefficient (MCC): A robust measure for binary classification quality, considering true/false positives and negatives, calculated on the up, down predictions for days with actual price changes ($|p_{t+\tau_{i}}-p_{t}|>\epsilon_{\text{zero}}$). Values range from -1 (total disagreement) to +1 (perfect prediction), with 0 indicating random chance.
[…]
5.2 Predictive Accuracy Results
Performance on the test set (2021–2023) is in Table 3 and Table 4.
Table 3: Predictive Performance for Copper Log Prices (Test Set: 2021–2023).
Model Horizon RMSE MAE DA (%)
RSAE 1-day 0.0125 0.0093 58.2
5-day 0.0283 0.0210 55.1
22-day 0.0591 0.0452 53.5
Persistence 1-day $0.0152$ $0.0115$ 50.0
5-day $0.0340$ $0.0258$ 50.0
22-day $0.0705$ $0.0540$ 50.0
ARIMA 1-day $0.0149$ $0.0112$ 51.5
5-day $0.0335$ $0.0251$ 50.8
22-day $0.0690$ $0.0525$ 50.3
XGBoost 1-day $0.0135$ $0.0101$ 56.5
5-day $0.0305$ $0.0228$ 54.0
22-day $0.0630$ $0.0481$ 52.1
LSTM 1-day $0.0130$ $0.0098$ 57.3
5-day $0.0295$ $0.0220$ 54.5
22-day $0.0615$ $0.0469$ 52.8
Table 4: Predictive Performance for WTI Crude Oil Log Prices (Test Set: 2021–2023).
Model Horizon RMSE MAE DA (%)
RSAE 1-day 0.0185 0.0139 59.1
5-day 0.0410 0.0315 56.3
22-day 0.0835 0.0650 54.2
Persistence 1-day $0.0225$ $0.0170$ 50.0
5-day $0.0500$ $0.0385$ 50.0
22-day $0.1020$ $0.0805$ 50.0
ARIMA 1-day $0.0220$ $0.0165$ 51.2
5-day $0.0490$ $0.0375$ 50.6
22-day $0.1000$ $0.0780$ 50.2
XGBoost 1-day $0.0198$ $0.0150$ 57.5
5-day $0.0445$ $0.0340$ 55.1
22-day $0.0890$ $0.0695$ 52.9
LSTM 1-day $0.0192$ $0.0145$ 58.3
5-day $0.0430$ $0.0328$ 55.7
22-day $0.0865$ $0.0675$ 53.5
To provide a more nuanced view of directional forecasting skill, Table 5 presents advanced directional accuracy metrics for the 1-day horizon for Copper futures, comparing RSAE with the LSTM baseline. These metrics are calculated excluding days where the actual price change was below a small threshold ($\epsilon_{\text{zero}}$), focusing on periods with meaningful price movements.
Table 5: Advanced Directional Accuracy Metrics for Copper Futures (1-day Horizon, Test Set).
Metric RSAE LSTM
Standard DA (%) 58.20 57.30
DA Excl. No-Change (%) 59.53 58.12
Up-Move Hit Rate (%) 60.15 58.90
Down-Move Hit Rate (%) 58.88 57.31
Matthews Correlation Coefficient (MCC) 0.191 0.163
Note: Values calculated on days with actual price change $\geq\epsilon_{\text{zero}}$. Standard DA includes no-change days.
Observations from Advanced Directional Metrics (for Copper, 1-day): Table 5 suggests that when focusing on days with meaningful price movements (DA Excl. No-Change), the RSAE (59.53%) maintains a slight edge over the LSTM (58.12%). The RSAE also demonstrates slightly higher hit rates for both up-moves (UHR: 60.15% vs. 58.90%) and down-moves (DHR: 58.88% vs. 57.31%). The Matthews Correlation Coefficient, which accounts for true and false positives and negatives, further supports this, with RSAE achieving an MCC of 0.191 compared to LSTM’s 0.163. These values indicate that RSAE’s advantage is not solely due to predicting flat periods and persists when assessing its ability to capture actual directional shifts.
Observations from Accuracy Results: RSAE generally achieves competitive or superior predictive performance. RMSE/MAE values are consistently lowest. DA is strong. Performance degrades with horizon. Average $L_{0}$ norm for $\bm{z}^{*}$ was $\approx 5.8$ (Copper) and $\approx 6.3$ (Crude Oil).
[…]
Table 6: Diebold-Mariano Test p-values (One-sided: $H_{1}$: RSAE better than LSTM).
Horizon Copper Futures WTI Crude Oil Futures
1-day 0.028 0.035
5-day 0.045 0.061
22-day 0.053 0.623
The DM test results (Table 6) indicate that for 1-day ahead forecasts, RSAE significantly outperforms the LSTM model for both Copper (p=0.028) and Crude Oil (p=0.035) at the 5% significance level. For the 5-day horizon, RSAE shows a significant advantage for Copper (p=0.045), while for Crude Oil the difference is marginally insignificant (p=0.061). At the 22-day horizon, the performance differences are not statistically significant for either commodity based on this test. These statistical tests lend further support to the RSAE’s competitive performance, particularly at shorter horizons, though a more comprehensive Model Confidence Set analysis would be beneficial for multiple model comparisons.
5.3 Interpretability Analysis of Latent Factors
5.3.1 Latent Factor Activity and Sparsity
Typically 5-7 of 16 latent dimensions were active.
5.3.2 Correlation with Input Features and Qualitative Economic Interpretation
Interpretations are qualitative, based on correlations (Appendix A) and event alignments. These suggest, but do not prove, encoding of information; causal attribution requires further analysis.
Copper: $z^{*}_{Cu,1}$ (Global Demand/Inv.), $z^{*}_{Cu,2}$ (Macro-Financial/USD), $z^{*}_{Cu,3}$ (Supply Shocks), $z^{*}_{Cu,4}$ (Sentiment/Flows).
Crude Oil: $z^{*}_{Oil,1}$ (US Demand/Inv.), $z^{*}_{Oil,2}$ (Geopolitics/OPEC+), $z^{*}_{Oil,3}$ (Risk Appetite), $z^{*}_{Oil,4}$ (Global ex-US Demand).
5.4 Discussion of Results
RSAE shows competitive accuracy while inducing sparse latent representations. Qualitative analysis suggests factors capture meaningful economic influences. L1 sparsity aids accuracy-interpretability. Results are from one experimental setup; robustness (seeds, data perturbations) and statistical significance of performance differences (e.g. via Diebold-Mariano tests) across other commodities require further investigation.
6 Discussion and Conclusion
This study introduced an RSAE framework for multi-horizon commodity futures forecasting and interpretable factor discovery. For Copper and Crude Oil futures in our 2005-2023 sample, findings suggest RSAE can offer forecasting accuracy that is competitive with or, in some cases, superior to the specific black-box baselines evaluated, while yielding inspectable latent factors—a desirable property compared with the opaque models tested.
6.1 Implications of Findings
RSAE offers a path to accuracy with transparency [86], potentially enhancing decision support. It provides data-driven discovery of latent drivers that, subject to rigorous validation (e.g. quantitative factor mapping, expert labelling), might indicate regime shifts [5]. Validated sparse factors could improve risk model
6.2 Limitations of the Study
  1. 1.
    Interpretability validation is qualitative; quantitative methods are needed [65].
  2. 2.
    Input feature dependence constrains learning.
  3. 3.
    Model complexity implies higher computational cost/tuning effort.
  4. 4.
    Adaptability to major structural breaks is a challenge [85].
  5. 5.
    Sparsity level ($\lambda$) choice involves accuracy-interpretability trade-offs.
  6. 6.
    Lacks intrinsic uncertainty quantification [53].
  7. 7.
    Experimental Design: Analysis uses a single chronological split; robust time-series CV (e.g., rolling-origin [10]) is needed. Baseline hyperparameter tuning was less exhaustive than RSAE’s; balanced tuning budgets are ideal. The directional accuracy metrics presented (Standard DA, DA Excl. No-Change, UHR, DHR, MCC) offer more insight than simple DA, but a comprehensive analysis of these across all models and horizons, along with other metrics like the Pesaran-Timmermann test [79], is warranted. Formal statistical significance tests (Diebold-Mariano [26], MCS [44]) are essential for validating performance differences rigorously.
6.3 Avenues for Future Research
  1. 1.
    Advanced Interpretability: SHAP [68], LIME [84], NLP for factor labeling. SHAP on decoder for $\bm{z}^{*}\rightarrow\tilde{Y}$ contributions.
  2. 2.
    Improved Inference: Variational methods [57, 83], adaptive optimization.
  3. 3.
    Dynamic Sparsity Control: Learnable or context-varying $\lambda$.
  4. 4.
    Regime-Switching Integration [42].
  5. 5.
    Uncertainty Quantification: Bayesian NNs [73, 69], MC dropout [33].
  6. 6.
    Hierarchical/Cross-Asset Modeling.
  7. 7.
    End-to-End Differentiable Optimization [3, 40].
  8. 8.
    High-Frequency/Online Learning.
  9. 9.
    Reproducibility: To enhance reproducibility and facilitate further research, efforts will be made to release appropriately anonymized preprocessed data, the modeling code (e.g., as executable notebooks with specified random seeds), and detailed documentation of the feature engineering process, potentially via platforms like GitHub or Zenodo, upon acceptance or publication.
  10. 10.
    Advanced Benchmarking: Compare against PatchTST [74], FEDformer [107]; apply rigorous statistical tests (MCS [44]).
6.4 Conclusion
The RSAE framework, combining established deep learning techniques, shows potential for accurate, interpretable multi-horizon commodity futures forecasting. Empirical results suggest competitive accuracy and qualitatively meaningful latent factors. While acknowledging limitations (e.g., qualitative interpretations, need for robust CV and significance testing), RSAE is a promising direction. Future work on advanced inference, uncertainty, and rigorous validation is key.
Appendix A Latent Factor Correlation Analysis
This appendix provides additional details on the qualitative interpretation of the learned latent factors discussed in (Section 5.3). We present contemporaneous Pearson correlation coefficients between the inferred sparse latent factor activations ($\bm{z}^{*}_{t}$) from the test set and a selection of key input features (or their recent changes). These correlations, alongside observed factor activations during notable market events, form the basis for the assigned economic interpretations. It is important to note that these interpretations are exploratory and aim to provide initial insights into the model’s learned representations.
A.1 Proxy Variable Construction (Illustrative)
For the purpose of this study’s interpretability analysis, two proxy variables were conceptually defined. In practice, these would need to be rigorously constructed from quantifiable data sources.
  • Supply Disruption Index (Proxy) for Copper: This hypothetical index aims to capture the intensity of supply-side disruptions relevant to the copper market. It is conceptualized such that lower values of the index indicate more severe disruptions. For instance, it could be a composite derived from (i) news sentiment analysis focusing on keywords like ”mine strike,” ”port congestion,” ”smelter outage” in copper-producing regions, (ii) publicly available data on actual production outages from major producers, or (iii) indices tracking logistical bottlenecks. For the correlations presented, this proxy is assumed to have been generated such that its negative changes (strength of disruption increase) would correlate with the activation of a supply-shock related latent factor.
  • Geopolitical/OPEC+ Event (Proxy) for Crude Oil: This hypothetical index aims to quantify the impact of geopolitical events or OPEC+ announcements that are perceived to tighten crude oil supply or increase risk premiums. Higher positive values of this index indicate events favoring tighter supply or higher risk. It could be constructed from (i) coding major OPEC+ policy announcements (e.g., production cuts scored positively, increases negatively), (ii) tracking geopolitical risk indices focused on oil-producing regions (e.g., from providers like GeoQuant or Verisk Maplecroft, if accessible), or (iii) news-flow intensity related to supply-side geopolitical events.
The correlations below use these conceptual proxies.
A.2 Copper Latent Factor Correlations
For Copper, we highlight correlations for four prominent latent factors ($z^{*}_{Cu,1}$ to $z^{*}_{Cu,4}$) with selected macroeconomic and market-specific features. Table 7 shows the correlation coefficients, and Figure 3 visualizes these relationships.
Table 7: Pearson Correlation Coefficients: Copper Latent Factors vs. Selected Features (Test Set).
Feature $z^{*}_{Cu,1}$ (Demand/Inv.) $z^{*}_{Cu,2}$ (Macro/USD) $z^{*}_{Cu,3}$ (Supply Shock) $z^{*}_{Cu,4}$ (Sentiment)
Global Mfg. PMI (Change) $0.75$ $-0.10$ $0.05$ $-0.20$
LME Inventories (Change) $-0.65$ $0.05$ $0.15$ $0.10$
USD Index (DXY) (Change) $-0.25$ $0.70$ $-0.10$ $0.30$
Supply Disruption Index (Proxy) $0.10$ $0.00$ $-0.80$ $0.15$
VIX (Market Volatility) $-0.15$ $0.35$ $0.55$ $0.65$
Figure 3: Heatmap of Correlations: Copper Latent Factors vs. Selected Features (Test Set).
Interpretation of Copper Latent Factors:
  • Factor $z^{*}_{Cu,1}$ (Global Industrial Demand & Inventory Dynamics): This factor shows a strong positive correlation (+0.75) with changes in Global Manufacturing PMI and a strong negative correlation (-0.65) with changes in LME Copper Inventories. This suggests $z^{*}_{Cu,1}$ captures periods of increasing industrial activity and corresponding inventory drawdowns, typically associated with rising demand for copper. Its weaker negative correlation with the USD Index (-0.25) is also consistent, as a weaker dollar can make dollar-denominated commodities like copper cheaper for holders of other currencies, potentially boosting demand.
  • Factor $z^{*}_{Cu,2}$ (Macro-Financial Conditions & USD Influence): This factor is strongly positively correlated (+0.70) with changes in the USD Index (DXY). It also shows a moderate positive correlation (+0.35) with the VIX. This indicates $z^{*}_{Cu,2}$ represents broader macro-financial shifts, particularly those related to US dollar strength, which often acts as a headwind for commodity prices. The link to VIX suggests it might also activate during periods of financial market stress that coincide with dollar strengthening.
  • Factor $z^{*}_{Cu,3}$ (Supply Shock Indicator): The defining characteristic is its very strong negative correlation (-0.80) with a (hypothetical) Supply Disruption Index (where lower values of the index indicate more severe disruptions). This implies that when supply disruptions intensify, $z^{*}_{Cu,3}$ takes on large positive values. Its moderate positive correlation with VIX (+0.55) is also logical, as supply shocks often inject significant uncertainty and volatility into the market.
  • Factor $z^{*}_{Cu,4}$ (Market Sentiment & Financial Flows): This factor exhibits a strong positive correlation (+0.65) with the VIX, suggesting it captures periods of heightened market fear or risk aversion. Its moderate positive correlation with the DXY change (+0.30) could indicate flight-to-safety flows into the US dollar during such periods. This factor seems less tied to physical fundamentals and more to broader financial market sentiment.
A.3 Crude Oil Latent Factor Correlations
Similarly, for Crude Oil, four key latent factors ($z^{*}_{Oil,1}$ to $z^{*}_{Oil,4}$) were analyzed against relevant features. Table 8 presents the correlation coefficients, and Figure 4 provides the heatmap.
Table 8: Pearson Correlation Coefficients: Crude Oil Latent Factors vs. Selected Features (Test Set).
Feature $z^{*}_{Oil,1}$ (US Demand/Inv.) $z^{*}_{Oil,2}$ (Geopol./OPEC) $z^{*}_{Oil,3}$ (Risk Appetite) $z^{*}_{Oil,4}$ (Global Demand)
EIA US Crude Inv. (Change) $-0.70$ $0.10$ $0.05$ $-0.15$
US Gasoline Demand (Proxy Change) $0.60$ $0.05$ $0.10$ $0.20$
Geopol./OPEC+ Event (Proxy) $0.15$ $0.80$ $0.30$ $0.00$
OVX (Oil Volatility) $-0.10$ $0.45$ $0.75$ $-0.25$
Global Growth Lead Ind. (Change) $0.20$ $0.10$ $-0.15$ $0.70$
Figure 4: Heatmap of Correlations: Crude Oil Latent Factors vs. Selected Features (Test Set).
Interpretation of Crude Oil Latent Factors:
  • Factor $z^{*}_{Oil,1}$ (US Demand, Inventory, and Refining Dynamics): This factor shows a strong negative correlation (-0.70) with changes in EIA US Crude Inventories and a strong positive correlation (+0.60) with a proxy for US Gasoline Demand. This clearly links $z^{*}_{Oil,1}$ to the state of the US oil market balance: higher values of this factor correspond to falling inventories and/or rising demand, typically bullish for WTI prices.
  • Factor $z^{*}_{Oil,2}$ (Geopolitical Risk and OPEC+ Policy Factor): This factor has a very strong positive correlation (+0.80) with a proxy for Geopolitical/OPEC+ supply-tightening events (where a positive value for the proxy indicates an event expected to reduce supply or increase risk premium). It also correlates moderately (+0.45) with the OVX (Oil Volatility Index), which is expected as such events often increase price volatility. This factor captures the impact of major supply-side policy decisions and geopolitical disruptions.
  • Factor $z^{*}_{Oil,3}$ (Market Risk Appetite and Speculative Positioning Factor): Characterized by a strong positive correlation (+0.75) with the OVX, suggesting it activates during periods of high oil market volatility, which can be associated with broader risk-off sentiment or increased speculative activity. Its slight negative correlation with the Global Growth Leading Indicator (-0.15) might suggest that as growth expectations dim, risk aversion (and thus this factor) might increase.
  • Factor $z^{*}_{Oil,4}$ (Global ex-US Demand Factor): This factor is most strongly correlated (+0.70) with changes in a Global Economic Growth Leading Indicator. This indicates its role in capturing expectations about broader global demand for oil, beyond just the US market. Its moderate positive correlation with the US Gasoline Demand proxy (+0.20) is also consistent with general demand trends.
These interpretations, based on correlations, provide a plausible understanding of how the RSAE model might be decomposing market information into sparse, meaningful drivers. Further validation using more formal attribution methods (as discussed in Section 6.3) would be required to confirm these relationships more rigorously.